Massive Stars as Cosmic Engines Proceedings IAU Symposium No. 250, 2008 F. Bresolin, P.A. Crowther & J. Puls, eds.

\begin{matrix}c \\ ◯ \end{matrix} 2008 International Astronomical Union

DOI: 00.0000/X000000000000000X

Young Massive Clusters

Donald F. Figer^{1}

1Chester F. Carlson Center for Imaging Science Rochester Institute of Technology Rochester, NY 14623-5604

                             email: figer@cis.rit.edu

Abstract. Over the past ten years, there has been a revolution in our understanding of massive young stellar clusters in the Galaxy. As recently as ten years ago, there were no known examples having masses >10^{4}, yet now we know that there are at least a half dozen such clusters in the Galaxy. In all but one case, the masses of these clusters have been determined through infrared observations. Several had been identified as clusters long ago, but their massive natures only recently determined. Presumably, we are just scratching the surface, and we might look forward to having statistically significant samples of coeval massive stars at all important stages of stellar evolution in the next ten years. I review the efforts that have led to this dramatic turn of events and the growing sample of young massive clusters in the Galaxy.

Keywords. star clusters, massive stars

1. Introduction

Massive stellar clusters are the birthplaces of massive stars. They are also the places where many massive stars reside all of their lives, having little time to wander before exploding as supernovae. The astrophysical importance of massive clusters largely derives from their content of massive stars, objects that have extraordinary effects on their surroundings.

Indeed, massive stars are key ingredients and probes of astrophysical phenomena on all size and distance scales, from individual star formation sites, such as Orion, to the early Universe during the age of reionization when the first stars were born. As ingredients, they control the dynamical and chemical evolution of their local environs and individual galaxies through their influence on the energetics and composition of the interstellar medium. They likely play an important role in the early evolution of the first galaxies, and there is evidence that they are the progenitors of the most energetic explosions in the Universe, seen as GRBs. As probes, they define the upper limits of the star formation process and their presence may end further formation of nearby lower mass stars and planets. They are also prominent output products of galactic mergers, starburst galaxies, and AGN. In their end states, i.e. supernovae, they produce the vast array of heavy elements that eventually build the ingredients for life.

Despite the importance of massive stars, and the clusters in which they reside, no truly massive clusters were known to exist in the Galaxy before about ten years ago, when the Arches and Quintuplet clusters were identified as being at least that massive (Figer et al. 1999b). Since then, a number of efforts have led to the identification of about a half dozen more such massive clusters in the Galaxy.

We are on the cusp of a revolution in massive stellar cluster research, as the identified sample is likely the “tip of the iceberg.” In the next ten years, we can expect to identify perhaps a factor of ten more massive clusters in the Galaxy than are currently known. With this sample, we can expect to routinely address many of the long-pursued questions

Table 1. Properties of massive clusters in the Galaxy^{a}

Log(M) Radius  Log(ρ) Age Log(L) Log(Q) OB YSG RSG LBV WN WC
Cluster  M_{⊙}  pc M_{⊙} pc^{−3} Myr  L_{⊙}  s^{−1}
Westerlund 1^{b}  4.7  1.0  4.1 3−6  ...  6  4  2 16  8
RSGC2^{c}  4.6  2.7  2.7 3−6  ...  ... 0  0  26  0  0  0
RSGC1^{d}  4.5  1.3  3.5 3−6  ...  ... 1  1  14  0  0  0
Quintuplet^{e}  4.3  1.0  3.2 3−6  7.5  50.9 100 100  1  2  6 13
Arches^{f}  4.3  0.19  5.6 2−2.5  8.0  51.0 160  0  0  0  6  0
Center^{g}  4.3  0.23  5.6 3−7  7.3  50.5 100  0  4  1 18 12
NGC 3603^{h}  4.1  0.3  5.0 2−2.5  ...  ... 60  0  0  0  3  0
Trumpler 14^{i}  4.0  0.5  4.3  ...  ...  ... 31  ...  ...  ... ... ...
Westerlund 2^{j}  4.0  0.8  3.7 3−6  ...  ... ...  ...  ...  ...  2 ...
Cl 1806-20^{k}  3.8  0.8  3.5 3−6  ...  ... 5  0  ...  1  2  2

aAn ellipsis has been entered in cases where data are not reliable or available. “M” is the total cluster mass in all stars extrapolated down to a lower-mass cutoff of 1 M_{⊙}, assuming a Salpeter IMF slope and an upper mass cutoff of 120 M_{⊙} (unless otherwise noted) “Radius” gives the average projected separation from the centroid position. “ρ” is M divided by the volume. This is probably closer to the central density than the average density because the mass is for the whole cluster while the radius is the average projected radius. “Age” is the assumed age for the cluster. “Luminosity” gives the total measured luminosity for observed stars. “Q” is the estimated Lyman continuum flux emitted by the cluster. ^{b}Figer et al. (2006). ^{c}Davies et al. (2007). dFiger et al. (2006). ^{e}Figer et al. (1999b). ^{f} Mass estimates have been made based upon the number of stars having M_{initial}>20 M_{⊙} given in Figer et al. (1999b) and the mass function slope in Kim et al. (2006). The age, luminosity and ionizing flux are from Figer et al. (2002). ^{g}Krabbe et al. (1995). The mass, M has been estimated by assuming that a total 10^{3.5} stars have been formed. The age spans a range covering an initial starburst, followed by an exponential decay in the star formation rate. ^{h}Harayama (2007). ^{i}Harayama (2007). jHarayama (2007). ^{k}Figer et al. (2005).

in massive star research to determine, for example: 1) the most massive star that can form, 2) the binary frequency of massive stars, 3) the properties of massive stellar winds, e.g. clumping, 4) the evolutionary sequence of massive stars, and 5) the end states of massive stars. Indeed, Barbosa & Figer (2004) summarize many interesting questions that might be addressed with future studies of massive stars.

In this paper, I review the Galactic sample of massive stellar clusters (M>10^{4} M_{⊙}), the efforts that led to their discovery, and the identification of a stellar upper mass limit. Finally, I speculate on the hidden population of massive stellar clusters in the Galaxy that will likely be revealed in the next ten years.

2. The sample of massive young clusters

There are approximately ten massive young clusters known in the Galaxy, with three being located within the central 50 pc of the Galactic center. Table 1 gives the known properties of these clusters in a list that is ordered with decreasing mass. In some cases, the properties are poorly determined, in which case the entries are filled with an ellipsis. The most massive, Westerlund 1, has a mass of ≈50,000 M_{⊙}. The densest are in the Galactic center, with the Arches cluster having a density of ≈300,000 M_{⊙} pc^{−3}. While most are quite young, the RSG (red supergiant) are distinctly older. Presumably, the table reflects an observational bias in that younger clusters are probably more concentrated and thus easier to identify as clusters.

In the following, I describe some of the ten massive clusters in the table and their massive stellar content.

2.1. Westerlund 1

Westerlund 1 is the most massive young cluster identified at this time in the Galaxy (Clark et al. 2005; Negueruela & Clark 2005; Skinner et al. 2006; Groh et al. 2006; Crowther et al. 2006). Commensurately, it is host to the most evolved massive stars in

Figure 1. Radio emission from the GC region at 6 cm, adapted in Figure 1 by Cotera et al. (1999) from Yusef-Zadeh & Morris (1987). The star symbols represent the three massive clusters.

the Galaxy, including half the known population of yellow supergiants and more WR stars than in any other cluster in the Galaxy. The cluster contains a magnetar, a highly magnetized neutron star that may be descended from a particularly massive progenitor Muno et al. (2006). Oddly, the massive stellar content in Westerlund 1 has only been recently revealed, more than 40 years after its discovery.

2.2. Central cluster

The Central cluster resides in the central parsec of the Galaxy and contains many massive stars formed in the past 10 Myr (Becklin et al. 1978; Rieke, Telesco, & Harper 1978; Lebofsky, Rieke, & Tokunaga 1982; Forrest et al. 1987; Allen, Hyland, & Hillier 1990; Krabbe et al. 1991; Najarro et al. 1994; Krabbe et al. 1995; Najarro 1995; Blum, Sellgren, & Depoy 1995b; Genzel et al. 1996; Najarro et al. 1997). There are at least 80 massive stars in the Central cluster (Eisenhauer et al. 2005), including ≈50 OB stars on the main sequence and 30 more evolved massive stars (see Figure 2). These young stars appear to be confined to two disks (Genzel et al. 2003; Tanner et al. 2006). There is also a tight collection of a dozen or so B stars (the “s” stars) in the central arcsecond, highlighted in the small box in the figure. The formation of so many massive stars in the central parsec remains as much a mystery now as it was at the time of the first infrared

Figure 2. K-band image of the Central cluster obtained with NAOS/CONICA from Schödel et al. (2006). The 100 or so brightest stars in the image are evolved descendants from main sequence O-stars. The central box highlights the “s” stars that are presumably young and massive (M_{initial}≈20 M_{⊙}).

observations of the region. Most recently, this topic has largely been supplanted by the even more improbable notion that star formation can occur within a few thousand AU of the supermassive black hole. See Figer (2008), and references therein, for a review of massive star formation and the “s” stars in the Galactic center.

2.3. Arches cluster

The Arches cluster is the densest young cluster in the Galaxy, and it has a relatively young age (Figer et al. 2002). Being so young and massive, it contains the richest collection of O-stars and WNL stars in any cluster in the Galaxy (Cotera et al. 1996; Serabyn, Shupe, & Figer 1998; Figer et al. 1999b; Blum et al. 2001; Figer et al. 2002; Figer 2005). Given its unique characteristics, it is ideal as a testbed for measuring the upper mass cutoff to the IMF (see Section 3). The brightest members of the cluster are strong emissionline stars (Harris et al. 1994; Nagata et al. 1995; Cotera 1995; Figer 1995; Cotera et al. 1996; Figer et al. 1999b; Blum et al. 2001; Figer et al. 2002). These lines are produced in stellar winds that are also detected at radio wavelengths (Lang, Goss, & Rodríguez 2001; Yusef-Zadeh et al. 2003; Lang et al. 2005; Figer et al. 2002), and x-ray wavelengths (Yusef-Zadeh et al. 2002; Rockefeller et al. 2005; Wang, Dong, & Lang 2006).

2.4. Quintuplet cluster

The Quintuplet cluster was originally noted for its five very bright stars (Glass, Moneti, & Moorwood 1990; Okuda et al. 1990; Nagata et al. 1990), but is now known to contain many massive stars (Geballe et al. 1994; Figer, McLean, & Morris 1995; Timmermann et al. 1996; Figer et al. 1999a). It is ≈4 Myr old and had an initial mass of >10^{4} M_{⊙} (Figer et al. 1999a). The hot stars in the cluster ionize the nearby “Sickle” HII region

JD 11. Young Massive Clusters 5

(see Figure 1). The Quintuplet is most similar to Westerlund 1 in mass, age, and spectral content (Clark et al. 2005; Negueruela & Clark 2005; Skinner et al. 2006; Groh et al. 2006; Crowther et al. 2006). Some of the stars in the cluster have been detected at x-ray wavelengths (Law & Yusef-Zadeh 2004), and at radio wavelengths (Lang et al. 1999, 2005). Recently, Tuthill et al. (2006) convincingly show that the five red stars in the cluster are dusty WC stars, characteristic of binary systems containing WCL plus an OB star (Tuthill, Monnier, & Danchi 1999). This may indicate that either the binary fraction for massive stars is extremely high (Nelan et al. 2004), or only binary massive stars evolve through the WCL phase (van der Hucht 2001). The Quintuplet cluster also contains two Luminous Blue Variables, the Pistol star (Harris et al. 1994; Figer et al. 1998, 1999c), and FMM362 (Figer et al. 1999a; Geballe, Najarro, & Figer 2000). Both stars are extraordinarily luminous (L>10^{6} L_{⊙}), yet relatively cool (T≈10^{4} K), placing them in the “forbidden zone” of the Hertzsprung-Russell Diagram, above the Humphreys-Davidson limit (?). The Pistol star is particularly intriguing, in that it is surrounded by one of the most massive (10 M_{⊙}) circumstellar ejecta in the Galaxy (Figer et al. 1999c; Smith 2006). Both stars are spectroscopically (Figer et al. 1999a) and photometrically variable (Glass et al. 2001).

2.5. Red Supergiant clusters

Figer et al. (2006), Davies et al. (2007), and Davies et al. (2008), identified two massive clusters in the red supergiant (RSG) phase, the first containing 14 (see Figure 3) and the second containing 26 RSGs, or collectively about 20% of all such stars known in the Galaxy. The inferred cluster masses are large, ≈3(10^{4}) M_{⊙} for the former and ≈4(10^{4}) M_{⊙} for the latter. The first cluster was first identified as a candidate in the Bica et al. (2003) catalog, and some of the stars the second cluster had already been identified as RSGs in Stephenson (1990) and Nakaya et al. (2001). Just like in the case of Westerlund 1, we see that a cluster once thought to be of relatively low mass can turn out to be quite massive on further inspection. Perhaps there are more massive clusters amongst the many clusters that have already been identified.

3. An upper limit to the masses of stars

Massive star clusters can be useful testing grounds for a variety of theoretical predictions, e.g. the IMF, binary fraction, n-body interactions, etc. Although only recently discovered, the currently known set of Galactic massive clusters have already yielded an interesting result regarding the upper limit to the masses of stars.

Theoretically, one might expect stellar mass to be limited by pulsational instabilities (Schwarzschild Härm 1959) or radiation pressure (Wolfire & Cassinelli 1987). Although stellar evolution models have been computed for massive stars up masses of 1000 M_{⊙}, no such stars have ever been observed. Indeed, some of the most famous “massive stars” have turned out to be multiple systems.

Observationally, the problem is difficult because massive stars are rare. They are formed in small numbers with respect to lower mass stars and they only live a few million years. A star formation event must produce about 10^{4} M_{⊙} in stars to have a statistically meaning expectation of stars with masses greater than 150 M_{⊙}. For example, such a cluster should have about three such stars, assuming a Salpeter initial mass function (Salpeter 1955) extrapolated to zero probability. Unfortunately, the cluster must satisfy a number of other criteria in order to use it for identifying an upper mass cutoff. For instance, it must be young enough so that its most massive members have not yet exploded as supernovae. Yet, it must be old enough to be free of natal molecular material. In order to identify its

Figure 3. GLIMPSE image of RSGC1. W42, to the upper left, is an unrelated star formation region along the line of sight.

individual stars, it must be close enough to us. To relate apparent magnitude to absolute magnitude, it must be at known distance. The stars must be coeval enough so that the star formation episode that produced the cluster can be considered to constitute a single event. Finally, the age must be known relatively accurately.

It is difficult to satisfy all these criteria, and, indeed, only the Arches cluster does. Being in the Galactic center happens to be useful in this regard, as its distance is then relatively well known as compared to other clusters. The result for the Arches cluster is shown in Figure 4. In this plot, it is apparent that there is an absence of stars with masses above about 150 M_{⊙}, where many are expected, i.e. there is an upper mass cutoff. See Weidner & Kroupa (2004), Weidner & Kroupa (2006) and Oey & Clarke (2005) for further arguments for such a cutoff.

4. The hidden population of massive stellar clusters

2MASS (Skrutskie et al. 1997) and GLIMPSE (Benjamin et al. 2003) have heralded a new era in massive star cluster research. With these surveys, we will be able to probe much further into the Galactic plane than ever before. We can expect an order of magnitude increase in the number of known young clusters in the Galaxy as these surveys are further investigated. Figure 5 shows the Galactic distribution of known young clusters from WEBDA (dots), candidate clusters from Bica et al. (2003) (triangles), and the verified massive clusters discussed in this paper, with the addition of several very likely massive clusters (octogons). The Galactic center is at (0,0) and the Sun is at (0,8). The visual clusters are mostly within 3 kpc from the Sun. The infrared clusters are a bit further away; however, it is clear that the number is highly incomplete and that the far side of the Galaxy as well as the central regions must contain many new clusters yet

Figure 4. Number versus mass for stars in the Arches cluster from Figer (2005). There is a clear deficit of stars with initial masses greater than 150 M_{⊙}, as seen in the hatched regions, for a reasonable range of IMF slopes.

to be discovered. We estimate that the total number of Galactic stellar clusters should exceed 20,000.

Acknowledgements

I thank the following individuals for discussions related to this work: Paco Najarro, Rolf Kudritzki, Maria Messineo, Ben Davies, Lucy Hadfield, Qingfeng Zhu, and Sungsoo Kim. The material in this paper is based upon work supported by NASA under award No. NNG05-GC37G, through the Long Term Space Astrophysics program. This research has made use of the SIMBAD database, Aladin and IDL software packages, and the GSFC IDL library. This research was performed in the Rochester Imaging Detector Laboratory with support from a NYSTAR Faculty Development Program grant. Some of the data presented herein were obtained at the W. M. Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation.

References

Alexander, T. 1999 The Distribution of Stars near the Supermassive Black Hole in the Galactic

Center. Astrophysical Journal 527, 835 Allen, D. A., Hyland, A. R., & Hillier, D. J. 1990 The source of luminosity at the Galactic

Centre. Monthly Notices of the Royal Astronomical Society 244, 706 Barbosa, C., & Figer, D. 2004 Top 10 Problems on Massive Stars. ArXiv Astrophysics e-

prints, arXiv:astro-ph/0408491 Becklin, E. E., Matthews, K., Neugebauer, G., & Willner, S. P. 1978 Infrared observa-

8 Donald F. Figer

tions of the galactic center. I - Nature of the compact sources. Astrophysical Journal 219, Figure 5. Galactic distribution of known young clusters (dots), candidate clusters (triangles), and verified massive clusters (octogons). The Galactic center is at (0,0) and the Sun is at (0,8). Distance for the known young clusters are from Dias et al. (2002). Distances for the candidate clusters were determined by fitting the line-of-sight velocities of nearby HII regions (Kuchar & Bania 1994) to the Galactic rotation curve, assuming that these regions are associated with the clusters. This figure is courtesy of Maria Messineo.

Benjamin, R. A., et al. 2003 GLIMPSE. I. An SIRTF Legacy Project to Map the Inner

Galaxy. Publications of the Astronomical Society of the Pacific 115, 953 Bica, E., Dutra, C. M., Soares, J., & Barbuy, B. 2003 New infrared star clusters in the

Northern and Equatorial Milky Way with 2MASS. Astronomy & Astrophysics 404, 223 Blum, R. D., Schaerer, D., Pasquali, A., Heydari-Malayeri, M., Conti, P. S., &

Schmutz, W. 2001 2 Micron Narrowband Adaptive Optics Imaging in the Arches Cluster. Astronomical Journal 122, 1875

Blum, R. D., Sellgren, K., & Depoy, D. L. 1995b Discovery of a possible Wolf-Rayet star

at the galactic center. Astrophysical Journal Letters 440, L17 Blum, R. D., Sellgren, K., & Depoy, D. L. 1996a Really Cool Stars at the Galactic Center.

Astronomical Journal 112, 1988 Clark, J. S., Negueruela, I., Crowther, P. A., & Goodwin, S. P. 2005 On the massive

stellar population of the super star cluster Westerlund 1. Astronomy & Astrophysics 434, 949

Cotera, A. S. 1995 Stellar Ionization of the Thermal Emission Regions in the Galactic Center.

Ph.D. Thesis Cotera, A. S., Erickson, E. F., Colgan, S. W. J., Simpson, J. P., Allen, D. A., &

Burton, M. G. 1996 The Discovery of Hot Stars near the Galactic Center Thermal Radio Filaments. Astrophysical Journal 461, 750

Cotera, A. S., Simpson, J. P., Erickson, E. F., Colgan, S. W. J., Burton, M. G., & Allen, D. A. 1999 Isolated Hot Stars in the Galactic Center Vicinity. Astrophysical Journal 510, 747

Crowther, P.A., Hadfield, L. J., Clark, J. S., Negueruela, I., & Vacca, W. D. 2006 A census of the Wolf-Rayet content in Westerlund 1 from near-infrared imaging and spectroscopy. ArXiv Astrophysics e-prints, arXiv:astro-ph/0608356

JD 11. Young Massive Clusters 9

Davies, B., Figer, D. F., Kudritzki, R.-P., MacKenty, J., Najarro, F., & Herrero, A. 2007 A Massive Cluster of Red Supergiants at the Base of the Scutum-Crux Arm. Astrophysical Journal 671, 781

Davies, B., Figer, D. F., Law, C. J., Kudritzki, R.-P., Najarro, F., Herrero, A., & MacKenty, J. W. 2007 The cool supergiant population of the massive young star cluster RSGC1. ArXiv e-prints 711, arXiv:0711.4757

Dias, W. S., Alessi, B. S., Moitinho, A., & Lépine, J. R. D. 2002 New catalogue of optically

visible open clusters and candidates. Astronomy & Astrophysics 389, 871 Dutra, C. M., Bica, E., Soares, J., & Barbuy, B. 2003 New infrared star clusters in the

southern Milky Way with 2MASS. Astronomy & Astrophysics 400, 533 Eckart, A., Ott, T., & Genzel, R. 1999, The Sgr A* stellar cluster: New NIR imaging and

spectroscopy. Astronomy & Astrophysics 352, L22 Eisenhauer, F., et al. 2005 SINFONI in the Galactic Center: Young Stars and Infrared Flares

in the Central Light-Month. Astrophysical Journal 628, 246 Figer, D. F. 1995 A Search for Emission-Line Stars Near the Galactic Center. Ph.D. Thesis. Figer, D. F. 2003, Massive stars and the creation of our Galactic Center. IAU Symposium,

212, 487 Figer, D. F. 2004, Young Massive Clusters in the Galactic Center. ASP Conf. Ser. 322: The

Formation and Evolution of Massive Young Star Clusters, 322, 49. Figer, D. F. 2005 An upper limit to the masses of stars. Nature 434, 192 Figer, D. F. 2008 Massive Star Formation in the Galactic Center. Massive Stars: From Pop III

and GRBs to the Milky Way, in press Figer, D. F., et al. 2000, 2 Micron Spectroscopy within 0.3” of Sagittarius A*. Astrophysical

Journal Letters 533, L49 Figer, D. F. et al. 2002 Massive Stars in the Arches Cluster. Astrophysical Journal 581, 258 Figer, D. F., et al. 2003 High-Precision Stellar Radial Velocities in the Galactic Center.

Astrophysical Journal 599, 1139 Figer, D. F., Becklin, E. E., McLean, I. S., & Morris, M. 1994 Discovery of Luminous NIR

Sources Associated With Ionized Gas Near the Galactic Center. ASSL Vol. 190: Astronomy with Arrays, The Next Generation, 545

Figer, D. F., Kim, S. S., Morris, M., Serabyn, E., Rich, R. M., & McLean, I. S. 1999b HST/NICMOS Observations of Massive Stellar Clusters Near the Galactic Center. Astrophysical Journal 525, 750.

Figer, D. F., MacKenty, J. W., Robberto, M., Smith, K., Najarro, F., Kudritzki, R. P., & Herrero, A. 2006 Discovery of an Extraordinarily Massive Cluster of Red Supergiants. Astrophysical Journal 643, 1166

Figer, D. F., McLean, I. S., & Morris, M. 1995 Two New Wolf-Rayet Stars and a Luminous Blue Variable Star in the Quintuplet (AFGL 2004) near the Galactic Center. Astrophysical Journal Letters 447, L29

Figer, D. F., McLean, I. S., & Morris, M. 1999a, Massive Stars in the Quintuplet Cluster.

Astrophysical Journal 514, 202 Figer, D. F., Morris, M., Geballe, T. R., Rich, R. M., Serabyn, E., McLean, I. S.,

Puetter, R. C., & Yahil, A. 1999c High-Resolution Infrared Imaging and Spectroscopy of the Pistol Nebula: Evidence for Ejection. Astrophysical Journal 525, 759

Figer, D. F., Najarro, F., Geballe, T. R., Blum, R. D., & Kudritzki, R. P. 2005 Massive

Stars in the SGR 1806-20 Cluster. Astrophysical Journal Letters 622, L49 Figer, D. F., Najarro, F., Morris, M., McLean, I. S., Geballe, T. R., Ghez, A. M., &

Langer, N. 1998 The Pistol star. Astrophysical Journal 506, 384 Figer, D. F., Rich, R. M., Kim, S. S., Morris, M., & Serabyn, E. 2004 An Extended

Star Formation History for the Galactic Center from Hubble Space Telescope NICMOS Observations. Astrophysical Journal 601, 319

Forrest, W. J., Shure, M. A., Pipher, J. L., & Woodward, C. E. 1987 Brackett Alpha

Images. AIP Conf. Proc. 155: The Galactic Center 155, 153 Frogel, J. A., Tiede, G. P., & Kuchinski, L. E. 1999 The Metallicity and Reddening of

Stars in the Inner Galactic Bulge. Astronomical Journal 117, 2296 Garay, G., & Lizano, S. 1999 Massive Stars: Their Environment and Formation. Publications

of the Astronomical Society of the Pacific 111, 1049 Geballe, T. R., Genzel, R., Krabbe, A., Krenz, T., & Lutz, D. 1994 Spectra of a Remark-

able Class of Hot Stars in the Galactic Center. ASSL Vol. 190: Astronomy with Arrays, The Next Generation, 73

Geballe, T. R., Najarro, F., & Figer, D. F. 2000 A Second Luminous Blue Variable in the

Quintuplet Cluster. Astrophysical Journal Letters 530, L97 Genzel, R., et al. 2003 The Stellar Cusp around the Supermassive Black Hole in the Galactic

Center. Astrophysical Journal 594, 812 Genzel, R., Eckart, A., Ott, T., & Eisenhauer, F. 1997, On the nature of the dark mass

in the centre of the Milky Way. Monthly Notices of the Royal Astronomical Society 291, 219

Genzel, R., Hollenbach, D., & Townes, C. H. 1994, The nucleus of our Galaxy. Reports of

Progress in Physics, 57, 417. Genzel, R., Thatte, N., Krabbe, A., Kroker, H., & Tacconi-Garman, L. E. 1996 The

Dark Mass Concentration in the Central Parsec of the Milky Way. Astrophysical Journal 472, 153

Genzel, R., & Townes, C. H. 1987 Physical conditions, dynamics, and mass distribution in

the center of the Galaxy. Annual Review of Astronomy and Astrophysics 25, 377 Ghez, A. M., Salim, S., Hornstein, S. D., Tanner, A., Lu, J. R., Morris, M., Beck-

lin, E. E., & Duchêne, G. 2005 Stellar Orbits around the Galactic Center Black Hole. Astrophysical Journal 620, 744

Glass, I. S., Matsumoto, S., Carter, B. S., & Sekiguchi, K. 2001 Large-amplitude variables near the Galactic Centre. Monthly Notices of the Royal Astronomical Society 321, 77

Glass, I. S., Moneti, A., & Moorwood, A. F. M. 1990 Infrared images and photometry of the cluster near G 0.15 - 0.05. Monthly Notices of the Royal Astronomical Society 242, 55P

Groh, J. H., Damineli, A., Teodoro, M., & Barbosa, C. L. 2006 Detection of additional Wolf-Rayet stars in the starburst cluster Westerlund 1 with SOAR. ArXiv Astrophysics e-prints, arXiv:astro-ph/0606498

Harayama, Y. 2007 The IMF of the massive star-forming region NGC 3603 from NIR adaptive

optics observations . PhD Thesis, LMU Mnchen Harris, A. I., Krenz, T., Genzel, R., Krabbe, A., Lutz, D., Politsch, A., Townes,

C. H., & Geballe, T. R. 1994 Spectroscopy of the Galactic Center Arches Region: Evidence for Massive Star Formation. NATO ASIC Proc. 445: The Nuclei of Normal Galaxies: Lessons from the Galactic Center, 223

Homeier, N. L., & Alves, J. 2005 Massive star formation in the W49 giant molecular cloud: Implications for the formation of massive star clusters. Astronomy & Astrophysics 430, 481

Homeier, N. L., Blum, R. D., Pasquali, A., Conti, P. S., & Damineli, A. 2003 Results from a near infrared search for emission-line stars in the Inner Galaxy: Spectra of new Wolf-Rayet stars. Astronomy & Astrophysics 408, 153

Kim, S. S., Figer, D. F., Lee, H. M., & Morris, M. 2000 N-Body Simulations of Compact

Young Clusters near the Galactic Center. Astrophysical Journal 545, 301 Kim, S. S., Figer, D. F., & Morris, M. 2004 Dynamical Friction on Galactic Center Star

Clusters with an Intermediate-Mass Black Hole. Astrophysical Journal Letters 607, L123 Kim, S. S., Figer, D. F., Najarro, F. N., & Kudritzki, R. P. 2006 Keck/LGS Observations

of the Arches Cluster. Astrophysical Journal in preparation Kim, S. S., & Morris, M. 2003 Dynamical Friction on Star Clusters near the Galactic Center.

Astrophysical Journal 597, 312 Koen, C. 2006 On the upper limit on stellar masses in the Large Magellanic Cloud cluster

R136. Monthly Notices of the Royal Astronomical Society 365, 590 Krabbe, A., et al. 1995 The Nuclear Cluster of the Milky Way: Star Formation and Velocity

Dispersion in the Central 0.5 Parsec. Astrophysical Journal Letters 447, L95 Krabbe, A., Genzel, R., Drapatz, S., & Rotaciuc, V. 1991 A cluster of He I emission-line

stars in the Galactic center. Astrophysical Journal Letters 382, L19 Kuchar, T. A., & Bania, T. M. 1994 Kinematic distances of Galactic H II regions from H I

absorption studies. Astrophysical Journal 436, 117

JD 11. Young Massive Clusters 11

Lang, C. C., Figer, D. F., Goss, W. M., & Morris, M. 1999 Radio Detections of Stellar Winds from the Pistol star and Other Stars in the Galactic Center Quintuplet Cluster. Astronomical Journal 118, 2327

Lang, C. C., Goss, W. M., & Rodríguez, L. F. 2001 Very Large Array Detection of the Ionized Stellar Winds Arising from Massive Stars in the Galactic Center Arches Cluster. Astrophysical Journal Letters 551, L143

Lang, C. C., Goss, W. M., & Wood, D. O. S. 1997 VLA H92 alpha and H115 beta Recombination Line Observations of the Galactic Center H II Regions: The Sickle (G0.18-0.04) and the Pistol (G0.15-0.05). Astrophysical Journal 474, 275

Lang, C. C., Johnson, K. E., Goss, W. M., & Rodríguez, L. F. 2005 Stellar Winds and Embedded Star Formation in the Galactic Center Quintuplet and Arches Clusters: Multifrequency Radio Observations. Astronomical Journal 130, 2185

Law, C., & Yusef-Zadeh, F. 2004 X-Ray Observations of Stellar Clusters Near the Galactic

Center. Astrophysical Journal 611, 858 Lebofsky, M. J., Rieke, G. H., & Tokunaga, A. T. 1982 M supergiants and star formation

at the galactic center. Astrophysical Journal 263, 736 Massey, P., & Hunter, D. A. 1998 Star Formation in R136: A Cluster of O3 Stars Revealed

by Hubble Space Telescope Spectroscopy. Astrophysical Journal 493, 180 Mercer, E. P., et al. 2005 New Star Clusters Discovered in the GLIMPSE Survey. Astro-

physical Journal 635, 560 Moffat, A. F. J., et al. 2002 Galactic Starburst NGC 3603 from X-Rays to Radio. Astro-

physical Journal 573, 191 Muno, M. P., et al. 2006 A Neutron Star with a Massive Progenitor in Westerlund 1. Astro-

physical Journal Letters 636, L41 Nagata, T., Woodward, C. E., Shure, M., & Kobayashi, N. 1995 Object 17: Another

cluster of emission-line stars near the galactic center. Astronomical Journal 109, 1676 Nagata, T., Woodward, C. E., Shure, M., Pipher, J. L., & Okuda, H. 1990 AFGL 2004

- an infrared quintuplet near the Galactic center. Astrophysical Journal 351, 83 Najarro, F. 1995 Quantitative Optical and Infrared Spectroscopy of Extreme Luminous Blue

Supergiants. Ph.D. Thesis Najarro, F., Figer, D. F., Hillier, D. J., & Kudritzki, R. P. 2004 Metallicity in the

Galactic Center: The Arches Cluster. Astrophysical Journal Letters 611, L105 Najarro, F., Hillier, D. J., Kudritzki, R. P., Krabbe, A., Genzel, R., Lutz, D., Dra-

patz, S., & Geballe, T. R. 1994 The nature of the brightest galactic center HeI emission line star. Astronomy & Astrophysics 285, 573

Najarro, F., Krabbe, A., Genzel, R., Lutz, D., Kudritzki, R. P., & Hillier, D. J. 1997 Quantitative spectroscopy of the HeI cluster in the Galactic center.. Astronomy & Astrophysics 325, 700

Nakaya, H., Watanabe, M., Ando, M., Nagata, T., & Sato, S. 2001 A Highly Reddened

Star Cluster Embedded in the Galactic Plane. Astronomical Journal 122, 876 Negueruela, I., & Clark, J. S. 2005 Further Wolf-Rayet stars in the starburst cluster West-

erlund 1. Astronomy & Astrophysics 436, 541 Nelan, E. P., Walborn, N. R., Wallace, D. J., Moffat, A. F. J., Makidon, R. B., Gies,

D. R., & Panagia, N. 2004 Resolving OB Systems in the Carina Nebula with the Hubble Space Telescope Fine Guidance Sensor. Astronomical Journal 128, 323

Oey, M. S., & Clarke, C. J. 2005 Statistical Confirmation of a Stellar Upper Mass Limit.

Astrophysical Journal Letters 620, L43 Okuda, H., et al. 1990 An infrared quintuplet near the Galactic center. Astrophysical Journal

351, 89 Portegies Zwart, S. F., Makino, J., McMillan, S. L. W., & Hut, P. 2002 The Lives and

Deaths of Star Clusters near the Galactic Center. Astrophysical Journal 565, 265 Portegies Zwart, S. F., McMillan, S. L. W., & Gerhard, O. 2003 The Origin of IRS 16:

Dynamically Driven In-Spiral of a Dense Star Cluster to the Galactic Center? Astrophysical Journal 593, 352

Rieke, G. H., Telesco, C. M., & Harper, D. A. 1978 The infrared emission of the Galactic center. Astrophysical Journal 220, 556 Rockefeller, G., Fryer, C. L., Melia, F., & Wang, Q. D. 2005 Diffuse X-Rays from the

Arches and Quintuplet Clusters. Astrophysical Journal 623, 171 Salpeter, E. E. 1955 The Luminosity Function and Stellar Evolution. Astrophysical Journal

121, 161 Sato, F., Hasegawa, T., Whiteoak, J. B., & Miyawaki, R. 2000 Cloud Collision-induced

Star Formation in Sagittarius B2. I. Large-Scale Kinematics. Astrophysical Journal 535, 857

Schwarzschild, M., Härm, R. 1959 On the Maximum Mass of Stable Stars. Astrophysical

Journal 129, 637 Schödel, R., et al. 2002 A star in a 15.2-year orbit around the supermassive black hole at

the centre of the Milky Way. Nature 419, 694 Schödel, R., Ott, T., Genzel, R., Eckart, A., Mouawad, N., & Alexander, T. 2003

Stellar Dynamics in the Central Arcsecond of Our Galaxy. Astrophysical Journal 596, 1015 Schödel, R., et al. 2006 Astronomy & Astrophysics , in preparation Serabyn, E., & Morris, M. 1996 Sustained star formation in the central stellar cluster of the

Milky Way.. Nature 382, 602 Serabyn, E., Shupe, D., & Figer, D. F. 1998 An extraordinary cluster of massive stars near

the centre of the Milky Way.. Nature 394, 448 Skinner, S. L., Simmons, A. E., Zhekov, S. A., Teodoro, M., Damineli, A., & Palla,

F. 2006 A Rich Population of X-Ray-emitting Wolf-Rayet Stars in the Galactic Starburst Cluster Westerlund 1. Astrophysical Journal Letters 639, L35

Skrutskie, M. F., et al. 1997 The Two Micron All Sky Survey (2MASS): Overview and

Status. The Impact of Large Scale Near-IR Sky Surveys 210, 25 Smith, N. 2006 Eruptive Mass Loss in Very Massive Stars and Population III Stars. ArXiv

Astrophysics e-prints, arXiv:astro-ph/0607457 Stephenson, C. B. 1990 A possible new and very remote galactic cluster. Astronomical Journal

99, 1867 Stolte, A. 2003, PhD Thesis, University of Heidelberg Stolte, A., Brandner, W., Grebel, E. K., Figer, D. F., Eisenhauer, F., Lenzen, R., &

Harayama, Y. 2003 NAOS-CONICA performance in a crowded field - the Arches cluster. The Messenger 111, 9

Stolte, A., Brandner, W., Grebel, E. K., Lenzen, R., & Lagrange, A.-M. 2005 The Arches Cluster: Evidence for a Truncated Mass Function? Astrophysical Journal Letters 628, L113

Stolte, A., Grebel, E. K., Brandner, W., & Figer, D. F. 2002 The mass function of the

Arches cluster from Gemini adaptive optics data. Astronomy & Astrophysics 394, 459 Tanner, A., et al. 2006 High Spectral Resolution Observations of the Massive Stars in the

Galactic Center. Astrophysical Journal 641, 891 Timmermann, R., Genzel, R., Poglitsch, A., Lutz, D., Madden, S. C., Nikola, T., Geis,

N., & Townes, C. H. 1996 Far-Infrared Observations of the Radio Arc (Thermal Arches) in the Galactic Center. Astrophysical Journal 466, 242

Tuthill, P. G., Monnier, J. D., & Danchi, W. C. 1999 A dusty pinwheel nebula around

the massive star WR 104.. Nature 398, 487 Tuthill, P., Monnier, J., Tanner, A., Figer, D., & Ghez, A. 2006 “Pinwheels” discovered

in the Quintuplet cluster. Science 313, 935. van der Hucht, K. A. 2001 The VIIth catalogue of galactic Wolf-Rayet stars. New Astronomy

Review 45, 135 Walborn, N. R., Maíz-Apellániz, J., & Barbá, R. H. 2002 Further Insights into the Struc-

ture of 30 Doradus from the Hubble Space Telescope Instruments. Astronomical Journal 124, 1601

Wang, Q. D., Dong, H., & Lang, C. 2006 The Interplay between Star Formation and the Nuclear Environment of our Galaxy: Deep X-ray Observations of the Galactic Center Arches and Quintuplet Clusters. ArXiv Astrophysics e-prints, arXiv:astro-ph/0606282

Weidner, C., & Kroupa, P. 2004 Evidence for a fundamental stellar upper mass limit from

clustered star formation. Monthly Notices of the Royal Astronomical Society 348, 187 Weidner, C., & Kroupa, P. 2006 The maximum stellar mass, star-cluster formation and

JD 11. Young Massive Clusters 13

composite stellar populations. Monthly Notices of the Royal Astronomical Society 365, 1333

Wolfire, M. G., & Cassinelli, J. P. 1987 Conditions for the formation of massive stars.

Astrophysical Journal 319, 850 Yusef-Zadeh, F., Law, C., Wardle, M., Wang, Q. D., Fruscione, A., Lang, C. C., &

Cotera, A. 2002 Detection of X-Ray Emission from the Arches Cluster near the Galactic Center. Astrophysical Journal 570, 665

Yusef-Zadeh, F., & Morris, M. 1987 Structural details of the Sagittarius A complex - Evidence for a large-scale poloidal magnetic field in the Galactic center region. Astrophysical Journal 320, 545

Yusef-Zadeh, F., Nord, M., Wardle, M., Law, C., Lang, C., & Lazio, T. J. W. 2003 Nonthermal Emission from the Arches Cluster (G0.121+0.017) and the Origin of gammaray Emission from 3EG J1746-2851. Astrophysical Journal Letters 590, L103

Zhao, J.-H., Desai, K., Goss, W. M., & Yusef-Zadeh, F. 1993 VLA Radio Recombination Line Observations of Ionized Gas in the -30 Kilometers per Second Molecular Cloud (G0.04+0.03) near the Galactic Center. I. The Discrete Radio Sources. Astrophysical Journal 418, 235

Discussion Vanbeveren: In 1982, I published a study where it was shown (from a theoretical point of view) that the maximum mass of the stars in a cluster depends on the total mass of the cluster and, from the cluster (total) mass function, I proposed a mass function for the maximum mass. This idea has been picked up by Weidner and Kroupa in 2004. Is there any observational evidence that he maximum mass of the stars in a cluster depends on the cluster mass?

Figer: Some work to answer that question has been done by Oey & Clarke (2005); however, the clusters in that study have very uncertain ages and it is thus not clear if their most massive members have already exploded as supernovae.

Damineli: Your open cluster candidates do not fit the ”radio” portion of the arms, specifically in the carina arm. How did you assign distances to the clusters?

Figer: We considered clusters associated with HII regions for which the line-of-velocities are available (Kuchar et al. 1994). We therefore obtained distances using the equations by Burton (1988) and assuming near-distances.

Langer: You use the argument that no WNE or WC stars are found in the Arches cluster to prove that it is young enough so that the most massive stars have not exploded off yet. Is this challenged by the evidence put forward by Nathan Smith that the most massive stars die as hydrogen-rich LBVs?

Figer: Nathan Smith’s speculation suggests that the most massive stars die as WNh. If this is true, then it is possible that we can no longer see stars with initial masses above 150 M_{⊙} in the Arches cluster because they have already exploded as supernovae, NOT because they did not form. This would also be true for R136 in 30 Dor. One interesting observational test of such a claim is to search for coeval clusters in which there are simultaneously WNh stars and more evolved subtypes (presumably from the highest initial masses in the cluster). Perhaps NGC 3603 is promising for such a test where there are WNh stars and a more evolved subtype, Sher 25; however, one would have to clearly show that these stars are coeval.